OpenStudy (anonymous):

Solve using the elimination method 3x+2y=3 and 9x-8y=-2

OpenStudy (anonymous):

I mean substitution method!!

OpenStudy (anonymous):

umm use Google.com yo

OpenStudy (anonymous):

plus yu dont know how to do like terms...woW

OpenStudy (anonymous):

x = 2/3 y = 1/2

OpenStudy (anonymous):

3x=(3-2y)

OpenStudy (anonymous):

9x=3(3x)